OpenStudy (anonymous):

Solve the following problem using substitution or elimination: To skate at Roller Heaven, each person must pay a monthly membership fee (m) and a fee for each session (f). In January, Art attended 5 sessions and paid $25 in all. In July, he attended 10 sessions, and he paid $40 in all. How much is the membership fee and how much are the session? Be sure to show all work.

OpenStudy (anonymous):

1 The monthly membership fee is a constant, therefore, no matter how many sessions Art goes to, that fee will stay the same we will denote this as (x) The fee for each session is a fee that is multiplied by the number of sessions he goes to, this will be denoted as (y) First equation: He did 5 sessions, therefore he paid the session fee five times, and the membership fee only once (Because it is a constant) so the equation is: 5y+x=25 Second Equation: Using the same logic as before, we get the second equation for the month of July: 10y+x=40 Now we will substitute. Solve the first equation for x in terms of y to get x=25-5y and plug this into the second equation 10y+25-5y=40 Now that we have the equation in terms of one variable, solve for y: 5y=15 y=3 The session fee is $3 Plug y into the first equation to get: 30+x=40 and you get x=10 the membership fee is $10

OpenStudy (anonymous):

Thanks Zhang

OpenStudy (anonymous):

Monthly fee + 5f = $25 Monthly fee + 10f = $40 (Monthly fee - Monthly fee) + (10f - 5f) = ($40 - $25) 5f = $15 f/5 = $15/5 Divide 15/5. Monthly fee + 5($3 sessions) = $25 Monthly fee + $15 = $25 - $15 - $15 Monthly fee = $10 and $3 per session.